OpenStudy (anonymous):

well, I know now how to find the (mean) average decrease but how about if I was trying to find the(mean) average increase?

OpenStudy (anonymous):

@iGreen would I just add them instead of subtracting?

OpenStudy (igreen):

Yes..do you have a specific question I can look at to confirm it?

OpenStudy (anonymous):

yes one sec.

OpenStudy (anonymous):

The school board makes budget decisions based on changes in enrollment. The board looked at the changes in enrollment at 3 middle schools. In school A enrollment increased from 4,320 students to 4,630. In school B enrollment increased from 4,820 to 5,740. In school C enrollment increased from 4,650 to 5,580. What is the mean (average) increase in the number of students for these three schools?

OpenStudy (igreen):

No you will subtract again.

OpenStudy (igreen):

You want to find the increase between these terms, so you will still subtract.

OpenStudy (anonymous):

ok, it's a good thing I didn't go with my own instict

OpenStudy (igreen):

Subtract terms: School A: 4,630 - 4320 = 310 Just think about it. You want to find the increase between these two numbers.. if you added both of them you will get 8950. And it certainly didn't increase by that much. School B: 5,740 - 4,820 = 920 School C: 5,580 - 4,650 = 930 Now that we have our three numbers(310, 920, 930) can you find the mean of them? @Blu-Girl

OpenStudy (anonymous):

3,830 right?

OpenStudy (anonymous):

actually no one sec

OpenStudy (igreen):

Lol.

OpenStudy (anonymous):

720

OpenStudy (lichking):

just a sec

OpenStudy (lichking):

aye 'Tis 720 so it is

OpenStudy (anonymous):

Thanks!

OpenStudy (lichking):

*bows*

OpenStudy (anonymous):

lol

OpenStudy (lichking):

*kneels*